Finding the area of a square,if mr.cruz want to construct a square table and has the area of 3 m^2,what shuold be the sulotion?
Answer by Boreal(15235)

A=3m^2
side is = s
A=s^2=3 m^2
s= sqrt(3)m=1.732 m as an approximation